OpenStudy (anonymous):

PLEASE HELP What is the radius of a circle with an area of 78.5 square inches?

OpenStudy (anonymous):

\[A = \pi(r)^2\] If this is an equation we could use, how would you go about plugging in values?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

im not very bright help meh

OpenStudy (anonymous):

\[78.5 = \pi (r)^2\] If you're solving for "r", and pi has a value of 3.14, don't we have everything except r?

OpenStudy (anonymous):

he stold that from me f you're solving for "r", and pi has a value of 3.14, don't we have everything except r?

OpenStudy (anonymous):

@youngthug1345 I've never met you before. ¯\_(ツ)_/¯

OpenStudy (anonymous):

yo cuhh trouvalii is my cuhh no one toches him but his mom

OpenStudy (anonymous):

Right, well... What you need to do to solve this equation is look for "r." Doing this requires you to move items around in your equation in order to get "r" alone. \[78.5 = \pi(r)^2\] would then equal \[\frac{ 78.5 }{ \pi } = r^2\] which would then turn into \[\sqrt{\frac{ 78.5 }{ \pi }} = r \].

OpenStudy (anonymous):

thank you @_Squish

OpenStudy (anonymous):

Sure thing.

OpenStudy (anonymous):

yo my new acount\